# Bayesian High-dimensional Linear Regression with Sparse Projection-posterior

### v2

@@ 4.2 Simulation Study 2: Distributed Computing and Credible Ellipsoid
This section shows the usefulness of the sparse projection-posterior method when the sample size is too huge. Besides the ability of our proposed method to accommodate distributed computing, the scope to jointly quantify for the signal uncertainty helps avoid the debiasing map. Since debiasing is computationally heavy, bypassing it gives even more computational edge over and above the distributed computation when $n$ is very large. We will conduct the following simulation to verify this claim. Here, we generate the design matrix from the normal distribution with $n=100000$ rows and $p=500$ columns. We set the sparsity to $s_{0}=20$ and generate the signals from the Uniform(1,3) distribution. We randomly divide the data into 100 splits to contain 1000 observations. Then, the $m$-th split of the data, containing the sub-matrix $\bm{X}_{m}\in\mathbb{R}^{1000\times 500}$ and the sub-vector $\bm{Y}_{m}\in\mathbb{R}^{1000}$ is sent to the $m$-th GPU to compute $\bm{X}_{m}^{\mathrm{T}}\bm{X}_{m}$ and $\bm{X}_{m}^{\mathrm{T}}\bm{Y}_{m}$, where $m=1,2,\dots,100$. Finally, these pieces are combined to get $\bm{X}^{\mathrm{T}}\bm{X}=\sum_{m=1}^{100}\bm{X}_{m}^{\mathrm{T}}\bm{X}_{m}$ and $\bm{X}^{\mathrm{T}}\bm{Y}=\sum_{m=1}^{100}\bm{X}_{m}^{\mathrm{T}}\bm{Y}_{m}$ following Section 2.5. Over 100 iterations, the average time taken for parallelly computing $\bm{X}_{m}^{\mathrm{T}}\bm{X}_{m}$ and $\bm{X}_{m}^{\mathrm{T}}\bm{Y}_{m}$ along with the cost of communication is only $0.17$ seconds. If, on the other hand, we used the entire data in a single machine to compute $\bm{X}^{\mathrm{T}}\bm{X}$ and $\bm{X}^{\mathrm{T}}\bm{X}$ directly, the average time taken over 100 replications is $18.32$ seconds. Although the rest of the method, including sampling from the vanilla posterior (10000 posterior samples with a burn-in of 1000) followed by the sparsity map, is carried out on a single machine, this initial distributed computing gives quite a computational edge to our method. Moreover, the horseshoe package in R that implements the horseshoe method cannot allocate the huge data size and hence is rendered incapable in such scenarios. We have not tried Bayesian LASSO or SSVS in this section as we already mentioned their high computation complexity in Section 4.1. The average estimation error (MSE) for the LASSO, sparse-projection, and sparse-projection distributed are respectively 0.0017, 0.00084, and 0.00097. The average MCC in the same order is 0.923, 0.998, and 0.981.

### v3

@@ 4.2 Simulation Study 2: Distributed Computing
This section shows the usefulness of the sparse projection-posterior method when the sample size is too huge. We will conduct the following simulation to verify this claim. Here, we generate the design matrix from the normal distribution with $n=100000$ rows and $p=500$ columns. We set the sparsity to $s_{0}=20$ and generate the signals from the Uniform(1,3) distribution. We randomly divide the data into 100 splits to contain 1000 observations. Then, the $m$-th split of the data, containing the sub-matrix $\bm{X}_{m}\in\mathbb{R}^{1000\times 500}$ and the sub-vector $\bm{Y}_{m}\in\mathbb{R}^{1000}$ is sent to the $m$-th GPU to compute $\bm{X}_{m}^{\mathrm{T}}\bm{X}_{m}$ and $\bm{X}_{m}^{\mathrm{T}}\bm{Y}_{m}$, where $m=1,2,\dots,100$. Finally, these pieces are combined to get $\bm{X}^{\mathrm{T}}\bm{X}=\sum_{m=1}^{100}\bm{X}_{m}^{\mathrm{T}}\bm{X}_{m}$ and $\bm{X}^{\mathrm{T}}\bm{Y}=\sum_{m=1}^{100}\bm{X}_{m}^{\mathrm{T}}\bm{Y}_{m}$ following Section 2.5. Over 100 iterations, the average time taken for parallelly computing $\bm{X}_{m}^{\mathrm{T}}\bm{X}_{m}$ and $\bm{X}_{m}^{\mathrm{T}}\bm{Y}_{m}$ along with the cost of communication is only $0.17$ seconds. If, on the other hand, we used the entire data in a single machine to compute $\bm{X}^{\mathrm{T}}\bm{X}$ and $\bm{X}^{\mathrm{T}}\bm{X}$ directly, the average time taken over 100 replications is $18.32$ seconds. Although the rest of the method, including sampling from the vanilla posterior (10000 posterior samples) followed by the sparsity map, is carried out on a single machine, this initial distributed computing gives quite a computational edge to our method. Moreover, the horseshoe package in R that implements the horseshoe method cannot allocate the huge data size and hence is rendered incapable in such scenarios. We have not tried Bayesian LASSO or SSVS in this section as we already mentioned their high computation complexity in Section 4.1. The average estimation error (MSE) for the LASSO, sparse-projection, and sparse-projection distributed are respectively 0.0017, 0.00084, and 0.00097. The average MCC in the same order is 0.923, 0.998, and 0.981.